OpenStudy (anonymous):

square root of ((a^6)(b^9))/-64

OpenStudy (oaktree):

So we take the square root of each individual part.\[\sqrt{\frac{ a^6*b^9 }{ -64 }} = \sqrt{\frac{ a^6*b^9 }{ 64 }}\sqrt{-1} = \frac{\left| a^3 \right| b^4i }{ 8 }\sqrt{b}\]

OpenStudy (anonymous):

yes I understand a little bit but im not understanding where the i and radical b came in